Eight uncharged capacitors with equal capacitances are combined in series. The combination is connected to a 6.53-V battery, which charges the capacitors. The charging process involves 3.77
Answer #1

Start with the formula Q = CV

3.77 X 10-4 = (C)(6.53)

C = 5.77 X 10-5 F

That is the equivalent capacitance of the system. Since the capacitors are in series, we need to apply the formula

1/CEQ = 1/C1 + 1/C2 + 1/C3 +...

Since all capacitors are equal, we get

1/CEQ = 8/C

1/(5.77 X 10-5) = 8/C

C = 4.62 X 10-4 F which can also be written as 462 μF or even .462 mF